# Infinitary combinatorics in condensed mathematics and strong homology

### v1

@@ 2.5. Consistent nonvanishing in higher degrees
First, this being the paper’s most combinatorially involved result, readers may wish to defer a reading of its proof until they’ve digested the argument of Theorem 4.10; its proof in fact will reference that argument at one point. That said, these references are essentially to motifs which readers familiar with $n$-coherence arguments will already know. Among the most central of these motifs is the following theorem [38]; this result is frequently instrumental, as in Theorems 2.20 and 4.10, in the construction of higher nontrivially coherent families.
@@
Let $I$ index an inverse system $\mathbf{X}$ of abelian groups. If the cofinality of $I$ is less than or equal to $\aleph_{k}$ then $\mathrm{lim}^{n}\,\mathbf{X}=0$ for all $n>k+1$.
If the transitive maps of $\mathbf{X}$ are all surjective, this vanishing holds even for all $n>k$, as observed in [82, Theorem 2.4]; in what follows, by Goblot’s Theorem we will mean both Theorem 2.15 and this corollary.
@@
If $Y=\kappa\times\lambda$, =$\mathcal{I}=\{X(f)\mid f\in{{}^{\kappa}}([\lambda]^{<\omega})\}$, and $\mathcal{J}=[Y]^{<\omega}$ then $\mathbf{X}[\mathcal{I},\mathcal{J},H]$ equals the system $\mathbf{A}_{\kappa,\lambda}[H]$ of Definition 2.4.
@@
$\mathrm{R}^{n}\mathrm{lim}_{I\in\mathcal{I}}\,\mathbf{X}[\mathcal{I},\mathcal{J},H]\neq 0$
@@
$\mathrm{R}^{n}\mathrm{lim}_{I\in\mathcal{I}}\,\mathbf{X}[\mathcal{I},\mathcal{J},H]\cong\mathrm{H}^{n}(\lambda;\mathcal{H})$
for all $n>0$, where the righthand terms are the sheaf cohomology groups of $\lambda$ (endowed with its order topology) with respect to the constant sheaf associated to $H$ (the argument is much as for [13, Thm. 3.2]).
@@
Let $\mu:=\aleph_{n}$, $H:=\bigoplus_{\omega_{n+1}}\mathbb{Z}$, and $\Lambda:={{}^{\mu}}([\mu^{+}]^{<\omega})$. We will begin by constructing an $(n+1)$-coherent family $\Phi^{\prime}=\langle\varphi^{\prime}_{\vec{\beta}}:\mu\times\bigwedge\vec{\beta}\rightarrow H^{\prime}\mid\vec{\beta}\in(\mu^{+})^{n+1}\rangle$, where $H^{\prime}$ will be either $\mathbb{Z}$ or $H$, depending on whether or not we are assuming the additional hypotheses in the statement of the theorem. Fix a sequence $\langle c_{\alpha}\mid\alpha<\mu^{+}\rangle$ witnessing that $S\in I[\mu^{+}]$. In particular, we may assume that each $c_{\alpha}$ is a subset of $\mu^{+}$ and, for all $\gamma\in S$, there is a club $d_{\gamma}$ in $\gamma$ such that
@@
  • for all $\beta<\gamma$, there is $\alpha<\gamma$ such that $d_{\gamma}\cap\beta=c_{\alpha}$.
@@
We now define $\varphi^{\prime}_{\vec{\beta}}$ for $\vec{\beta}\in(\mu^{+})^{n+1}$. Since we require $\Phi^{\prime}$ to be alternating, it suffices to define $\varphi^{\prime}_{\vec{\beta}}$ for strictly increasing sequences $\vec{\beta}=\langle\beta_{i}\mid i\leq n\rangle$. We do so by recursion on $\beta_{n}$. For terminological convenience, let “Case 1” denote the case in which we are not assuming any additional hypotheses and $H^{\prime}=\bigoplus_{\omega_{n+1}}\mathbb{Z}$, and let “Case 2” denote the case in which the additional hypotheses are being assumed and $H^{\prime}=\mathbb{Z}$. As part of our recursion hypotheses, we will require that, if we are in Case 1, then, for all increasing $\vec{\beta}\in(\mu^{+})^{n+1}$ and all $(\eta,\alpha)\in\mu\times\beta_{0}$, we have $\varphi^{\prime}_{\vec{\beta}}(\eta,\alpha)\in\bigoplus_{\beta_{n}+\mu}\mathbb{Z}$. If we are in Case 1, fix a nontrivial $n$-coherent family $\langle\tau_{\vec{\eta}}:\bigwedge\vec{\eta}\rightarrow\bigoplus_{\mu}\mathbb{Z}\mid\vec{\eta}\in(\omega_{n})^{n}\rangle$ (such exist by a slight modification of the argument of Theorem 4.10 below; alternately, see [12, Theorem 7.6]). Recall that, in Case 2, we already have a nontrivial $n$-coherent family $\langle\tau_{\vec{\eta}}:\bigwedge\vec{\eta}\rightarrow\mathbb{Z}\mid\vec{\eta}\in(\omega_{n})^{n}\rangle$. Given $\gamma<\mu^{+}$, we define a “shift homomorphism” $\mathrm{sh}_{\gamma}:\bigoplus_{\mu}\mathbb{Z}\rightarrow\bigoplus_{\gamma+\mu}\mathbb{Z}$ as follows: for all $x\in\bigoplus_{\mu}\mathbb{Z}$, let $\mathrm{sh}_{\gamma}(x)$ be the unique $y\in\bigoplus_{\gamma+\mu}\mathbb{Z}$ such that $\mathrm{supp}(y)=\{\gamma+\eta\mid\eta\in\mathrm{supp}(x)\}$ and, for all $\eta\in\mathrm{supp}(x)$, we have $y(\gamma+\eta)=x(\eta)$.
@@
Suppose now that we are in Case 2. We begin by defining auxiliary families $\bar{\Phi}^{\gamma}=\langle\bar{\varphi}_{\vec{\eta}}:\eta_{0}\rightarrow\mathbb{Z}\mid\vec{\eta}\in\mu^{n+1}\rangle$, $\bar{\Sigma}^{\gamma}=\langle\bar{\sigma}^{\gamma}_{\vec{\eta}}:\eta_{0}\rightarrow\mathbb{Z}\mid\vec{\eta}\in\mu^{n}\rangle$, and $\bar{\Psi}^{\gamma}=\langle\bar{\psi}^{\gamma}_{\vec{\eta}}:\eta_{0}\rightarrow\mathbb{Z}\mid\vec{\eta}\in\mu^{n}\rangle$. For the first, temporarily fix $\vec{\eta}\in\mu^{n+1}$, and let $\vec{\beta}\in\gamma^{n+1}$ be given by letting $\beta_{i}=\delta^{\gamma}_{\eta_{i}}$ for all $i\leq n$. Now, for all $\xi<\eta_{0}$, let $\bar{\varphi}_{\vec{\eta}}(\xi):=\varphi^{\prime}_{\vec{\beta}}(\xi,\delta^{\gamma}_{\xi})$. For the latter two, fix $\vec{\eta}\in\mu^{n}$, and let $\vec{\beta}\in\gamma^{n}$ be given by letting $\beta_{i}=\delta^{\gamma}_{\eta_{i}}$ for all $i<n$. First, for all $\xi<\eta_{0}$, let $\bar{\sigma}^{\gamma}_{\vec{\eta}}(\xi)=\sigma^{\gamma}_{\vec{\beta}}(\xi,\delta^{\gamma}_{\xi})$. Finally, to define $\bar{\psi}^{\gamma}_{\vec{\eta}}$, define $b\in\gamma^{n}$ by letting $b(i)$ be such that $d_{\gamma}\cap\beta_{i}=c_{b(i)}$ for all $i<n$, and let $\bar{\psi}^{\gamma}_{\vec{\eta}}(\xi)=\psi^{\gamma}_{b}(\delta^{\gamma}_{\xi})$ for all $\xi<\eta_{0}$.
By construction, $\bar{\Phi}^{\gamma}$ is $(n+1)$-coherent, and $\bar{\Sigma}^{\gamma}$ trivializes $\bar{\Phi}^{\gamma}$. If $\bar{\Psi}^{\gamma}$ does not trivialize $\bar{\Phi}^{\gamma}$, then proceed exactly as in the case in which $\gamma\notin S$, i.e., set $\varphi^{\prime}_{\vec{\beta}^{\frown}\langle\gamma\rangle}:=(-1)^{n}\sigma^{\gamma}_{\vec{\beta}}$ for all $\vec{\beta}\in\gamma^{n}$. Suppose now that $\bar{\Psi}^{\gamma}$ does trivialize $\bar{\Phi}^{\gamma}$. It follows that the family $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}=\langle\bar{\sigma}^{\gamma}_{\vec{\eta}}-\bar{\psi}^{\gamma}_{\vec{\eta}}\mid\vec{\eta}\in\mu^{n}\rangle$ is $n$-coherent. If $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is nontrivial, then again proceed as in the case in which $\gamma\notin S$. On the other hand, if $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is trivial, then proceed as follows. Suppose that $\vec{\beta}\in d^{n}_{\gamma}$, and let $\vec{\eta}\in\mu^{n}$ be such that $\beta_{i}=\delta^{\gamma}_{\eta_{i}}$ for all $i<n$. Then, for all $(\xi,\alpha)\in\mu\times\beta_{0}$, set
@@
It remains to show that $\Phi$ is nontrivial. Assume that $n>1$; the case $n=1$ is similar but easier. Suppose for sake of contradiction that $\Psi=\langle\psi_{\vec{f}}:X(\vec{f})\rightarrow H^{\prime}\mid\vec{f}\in\Lambda^{n}\rangle$ witnesses that $\Phi$ is trivial. For each $\alpha<\mu^{+}$, define a function $f_{\alpha}\in\Lambda$ by letting
@@
If $b\in(\mu^{+})^{\leq n}$ is such that, for all $i<j<|b|$, we have $c_{b(i)}\sqsubseteq c_{b(j)}$, then let $\vec{f}_{b}\in\Lambda^{|b|}$ be defined by letting $\vec{f}_{b}(i)=f_{b(i)}$ for all $i<|b|$, and, if $|b|=n$, then define $\psi^{*}_{b}:c_{b(0)}\rightarrow H^{\prime}$ by letting $\psi^{*}_{b}(c_{b(0)}(\eta))=\psi_{\vec{f}_{b}}(\eta,c_{b(0)}(\eta))$ for all $\eta<\mathrm{otp}(c_{b(0)})$. For all other $b\in(\mu^{+})^{n}$, simply let $\psi^{*}_{b}:\bigcap_{\alpha\in b}c_{\alpha}\rightarrow H^{\prime}$ be the constant function, taking value 00.
@@
Now assume that we are in Case 2, and fix a $\gamma\in S$ such that $A_{\gamma}=\langle\psi^{*}_{b}\mid b\in\gamma^{n}\rangle$. Define $g$ and $g_{\eta}$ for $\eta<\mu$ as in the previous case. Let $\bar{\Phi}^{\gamma}$, $\bar{\Sigma}^{\gamma}$, and $\bar{\Psi}^{\gamma}$ be as in stage $\gamma$ of the construction of $\Phi^{\prime}$. By our choice of $\gamma$, it follows that $\bar{\Psi}^{\gamma}$ trivializes $\bar{\Phi}^{\gamma}$. Suppose first that $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is nontrivial. Fix $\vec{\eta}\in\mu^{n}$, and let $b\in\gamma^{n}$ be such that $c_{b(i)}=\{\delta^{\gamma}_{\xi}\mid\xi<\eta_{i}\}$ for all $i<n$. Unraveling the definitions and constructions yields the following facts.
@@
    For all $\xi<\eta_{0}$, we have $\psi^{*}_{\vec{f}_{b}}(\delta^{\gamma}_{\xi})=\bar{\psi}^{\gamma}_{\vec{\eta}}(\xi)$.
@@
For $\vec{\zeta}\in\mu^{n-1}$, define $\nu_{\vec{\zeta}}:\zeta_{0}\rightarrow\mathbb{Z}$ as follows. Let $\vec{g}_{\vec{\zeta}}=\langle g_{\zeta_{i}}\mid i<n-1\rangle$, and let $\nu_{\vec{\zeta}}(\xi)=\psi_{\vec{g}_{\vec{\zeta}}{}^{\frown}\langle g\rangle}(\xi,\delta^{\gamma}_{\xi})$ for all $\xi<\zeta_{0}$. Then item (3) above implies that $\langle(-1)^{n}\nu_{\vec{\zeta}}\mid\vec{\zeta}\in\mu^{n-1}\rangle$ trivializes $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$, which is a contradiction.
Finally, assume that $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is trivial, as witnessed by $\langle\rho_{\vec{\zeta}}\mid\vec{\zeta}\in[\mu]^{n-1}\rangle$. Fix $\vec{\eta}\in\mu^{n}$ and $b\in\gamma^{n}$ as in the previous paragraph. Item (1) of that paragraph is unchanged, and items (2) and (3) become
@@
For all $\vec{\zeta}\in\mu^{n-1}$, define $\nu_{\vec{\zeta}}:\zeta_{0}\rightarrow\mathbb{Z}$ as follows. Let $\vec{g}_{\vec{\zeta}}=\langle g_{\zeta_{i}}\mid i<n-1\rangle$, and let $\nu_{\vec{\zeta}}(\xi)=(-1)^{n-1}\rho_{\vec{\zeta}}(\xi)+\psi_{\vec{g}_{\vec{\zeta}}{}^{\frown}\langle g\rangle}(\xi,\delta^{\gamma}_{\xi})$ for all $\xi<\zeta_{0}$. Then the above calculation implies that $\langle\nu_{\vec{\zeta}}\mid\vec{\zeta}\in\mu^{n-1}\rangle$ trivializes $\langle\tau_{\eta}\mid\vec{\eta}\in\mu^{n}\rangle$, yielding the final contradiction. ∎
@@
The case of $n=0$ is the ZFC result Theorem 2.10(5). For higher $n$, the relevant conditions of Theorem 2.20 all hold in Gödel’s constructible universe $\mathrm{L}$. To see this, first note that $\mathsf{GCH}$ holds in $\mathrm{L}$ and therefore, by [35, Theorem 3.40], $S^{\aleph_{n+1}}_{\aleph_{n}}\in I[\aleph_{n+1}]$ for all $1\leq n<\omega$. By a theorem of Jensen [47], $\diamondsuit(S^{\aleph_{n+1}}_{\aleph_{n}})$ holds in $\mathrm{L}$. Finally, there exists a nontrivial $n$-coherent family $\langle\tau_{\vec{\eta}}:\bigwedge\vec{\eta}\rightarrow\mathbb{Z}\mid\vec{\eta}\in(\omega_{n})^{n}\rangle$ in $\mathrm{L}$ by [15, Cor. 3.27]. ∎

### v2

@@ 2.5. Consistent nonvanishing in higher degrees
First, this being the paper’s most combinatorially involved result, readers may wish to defer a reading of its proof until they’ve digested the argument of Theorem 4.11; its proof in fact will reference that argument at one point. That said, these references are essentially to motifs which readers familiar with $n$-coherence arguments will already know. Among the most central of these motifs is the following theorem [41]; this result is frequently instrumental, as in Theorems 2.20 and 4.11, in the construction of higher nontrivially coherent families.
@@
Let $I$ be a directed partial order indexing an inverse system $\mathbf{X}$ of abelian groups. If $\mathrm{cf}(I)\leq\aleph_{k}$, then $\mathrm{lim}^{n}\,\mathbf{X}=0$ for all $n>k+1$.
If the transitive maps of $\mathbf{X}$ are all surjective, this vanishing holds even for all $n>k$, as observed in [85, Theorem 2.4]; in what follows, by Goblot’s Theorem we will mean both Theorem 2.15 and this corollary.
@@
[…]
@@
$\operatorname{lim}^{n}\mathbf{X}[\mathcal{I},\mathcal{J},H]\neq 0$
@@
$\operatorname{lim}^{n}\mathbf{X}[\mathcal{I},\mathcal{J},H]\cong\mathrm{H}^{n}(\lambda;\mathcal{H})$
for all $n>0$, where the righthand terms are the sheaf cohomology groups of $\lambda$ (endowed with its order topology) with respect to the constant sheaf associated to $H$ (the argument is much as for [14, Thm. 3.2]).
@@
[…]
@@
  • for all $\beta<\gamma$, there is an $\alpha<\gamma$ such that $d_{\gamma}\cap\beta=c_{\alpha}$.
@@
We now define $\varphi^{\prime}_{\vec{\beta}}$ for $\vec{\beta}\in(\mu^{+})^{n+1}$. Since we require $\Phi^{\prime}$ to be alternating, it suffices to define $\varphi^{\prime}_{\vec{\beta}}$ for strictly increasing sequences $\vec{\beta}=\langle\beta_{i}\mid i\leq n\rangle$. We do so by recursion on $\beta_{n}$. For terminological convenience, let “Case 1” denote the case in which we are not assuming any additional hypotheses and $H^{\prime}=\bigoplus_{\omega_{n+1}}\mathbb{Z}$, and let “Case 2” denote the case in which the additional hypotheses are being assumed and $H^{\prime}=\mathbb{Z}$. As part of our recursion hypotheses, we will require that, if we are in Case 1, then, for all increasing $\vec{\beta}\in(\mu^{+})^{n+1}$ and all $(\eta,\alpha)\in\mu\times\beta_{0}$, we have $\varphi^{\prime}_{\vec{\beta}}(\eta,\alpha)\in\bigoplus_{\beta_{n}+\mu}\mathbb{Z}$. If we are in Case 1, fix a nontrivial $n$-coherent family $\langle\tau_{\vec{\eta}}:\bigwedge\vec{\eta}\rightarrow\bigoplus_{\mu}\mathbb{Z}\mid\vec{\eta}\in(\omega_{n})^{n}\rangle$ (such exist by a slight modification of the argument of Theorem 4.11 below; alternatively, see [13, Theorem 7.6]). Recall that, in Case 2, we already have a nontrivial $n$-coherent family $\langle\tau_{\vec{\eta}}:\bigwedge\vec{\eta}\rightarrow\mathbb{Z}\mid\vec{\eta}\in(\omega_{n})^{n}\rangle$. Given $\gamma<\mu^{+}$, we define a “shift homomorphism” $\mathrm{sh}_{\gamma}:\bigoplus_{\mu}\mathbb{Z}\rightarrow\bigoplus_{\gamma+\mu}\mathbb{Z}$ as follows: for all $x\in\bigoplus_{\mu}\mathbb{Z}$, let $\mathrm{sh}_{\gamma}(x)$ be the unique $y\in\bigoplus_{\gamma+\mu}\mathbb{Z}$ such that $\mathrm{supp}(y)=\{\gamma+\eta\mid\eta\in\mathrm{supp}(x)\}$ and, for all $\eta\in\mathrm{supp}(x)$, we have $y(\gamma+\eta)=x(\eta)$.
@@
Suppose now that we are in Case 2. We begin by defining auxiliary families $\bar{\Phi}^{\gamma}=\langle\bar{\varphi}^{\gamma}_{\vec{\eta}}:\eta_{0}\rightarrow\mathbb{Z}\mid\vec{\eta}\in\mu^{n+1}\rangle$, $\bar{\Sigma}^{\gamma}=\langle\bar{\sigma}^{\gamma}_{\vec{\eta}}:\eta_{0}\rightarrow\mathbb{Z}\mid\vec{\eta}\in\mu^{n}\rangle$, and $\bar{\Psi}^{\gamma}=\langle\bar{\psi}^{\gamma}_{\vec{\eta}}:\eta_{0}\rightarrow\mathbb{Z}\mid\vec{\eta}\in\mu^{n}\rangle$. For the first, temporarily fix $\vec{\eta}\in\mu^{n+1}$, and let $\vec{\beta}\in\gamma^{n+1}$ be given by letting $\beta_{i}=\delta^{\gamma}_{\eta_{i}}$ for all $i\leq n$. Now, for all $\xi<\eta_{0}$, let $\bar{\varphi}^{\gamma}_{\vec{\eta}}(\xi):=\varphi^{\prime}_{\vec{\beta}}(\xi,\delta^{\gamma}_{\xi})$. For the latter two, fix $\vec{\eta}\in\mu^{n}$, and let $\vec{\beta}\in\gamma^{n}$ be given by letting $\beta_{i}=\delta^{\gamma}_{\eta_{i}}$ for all $i<n$. First, for all $\xi<\eta_{0}$, let $\bar{\sigma}^{\gamma}_{\vec{\eta}}(\xi)=\sigma^{\gamma}_{\vec{\beta}}(\xi,\delta^{\gamma}_{\xi})$. Finally, to define $\bar{\psi}^{\gamma}_{\vec{\eta}}$, define $b\in\gamma^{n}$ by letting $b(i)$ be such that $d_{\gamma}\cap\beta_{i}=c_{b(i)}$ for all $i<n$, and let $\bar{\psi}^{\gamma}_{\vec{\eta}}(\xi)=\psi^{\gamma}_{b}(\delta^{\gamma}_{\xi})$ for all $\xi<\eta_{0}$.
By construction, $\bar{\Phi}^{\gamma}$ is $(n+1)$-coherent, and $\bar{\Sigma}^{\gamma}$ trivializes $\bar{\Phi}^{\gamma}$. If $\bar{\Psi}^{\gamma}$ does not trivialize $\bar{\Phi}^{\gamma}$, then proceed exactly as in the case in which $\gamma\notin S$, i.e., set $\varphi^{\prime}_{\vec{\beta}^{\frown}\langle\gamma\rangle}:=(-1)^{n}\sigma^{\gamma}_{\vec{\beta}}$ for all $\vec{\beta}\in\gamma^{n}$. Suppose now that $\bar{\Psi}^{\gamma}$ does trivialize $\bar{\Phi}^{\gamma}$. It follows that the family $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}=\langle\bar{\sigma}^{\gamma}_{\vec{\eta}}-\bar{\psi}^{\gamma}_{\vec{\eta}}\mid\vec{\eta}\in\mu^{n}\rangle$ is $n$-coherent. If $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is nontrivial, then again proceed as in the case in which $\gamma\notin S$. On the other hand, if $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is trivial, then proceed as follows. Suppose that $\vec{\beta}\in(d_{\gamma})^{n}$, and let $\vec{\eta}\in\mu^{n}$ be such that $\beta_{i}=\delta^{\gamma}_{\eta_{i}}$ for all $i<n$. Then, for all $(\xi,\alpha)\in\mu\times\beta_{0}$, set
@@
It remains to show that $\Phi$ is nontrivial. Assume that $n>1$; the case $n=1$ is similar but easier. Suppose for the sake of contradiction that $\Psi=\langle\psi_{\vec{f}}:X(\vec{f})\rightarrow H^{\prime}\mid\vec{f}\in\Lambda^{n}\rangle$ witnesses that $\Phi$ is trivial. For each $\alpha<\mu^{+}$, define a function $f_{\alpha}\in\Lambda$ by letting
@@
If $b\in(\mu^{+})^{\leq n}$ is such that, for all $i<j<|b|$, we have $c_{b(i)}\sqsubseteq c_{b(j)}$, then let $\vec{f}_{b}\in\Lambda^{|b|}$ be defined by letting $\vec{f}_{b}(i)=f_{b(i)}$ for all $i<|b|$, and, if $|b|=n$, then define $\psi^{*}_{b}:c_{b(0)}\rightarrow H^{\prime}$ by letting $\psi^{*}_{b}(c_{b(0)}(\eta))=\psi_{\vec{f}_{b}}(\eta,c_{b(0)}(\eta))$ for all $\eta<\mathrm{otp}(c_{b(0)})$. For all other $b\in(\mu^{+})^{n}$, simply let $\psi^{*}_{b}:\bigcap_{\alpha\in b}c_{\alpha}\rightarrow H^{\prime}$ be the constant function, taking value 0.
@@
Now assume that we are in Case 2, and fix a $\gamma\in S$ such that $A_{\gamma}=\langle\psi^{*}_{b}\mid b\in\gamma^{n}\rangle$. Define $g$ and $g_{\eta}$ for $\eta<\mu$ as in the previous case. Let $\bar{\Phi}^{\gamma}$, $\bar{\Sigma}^{\gamma}$, and $\bar{\Psi}^{\gamma}$ be as in stage $\gamma$ of the construction of $\Phi^{\prime}$. By our choice of $\gamma$, it follows that $\bar{\Psi}^{\gamma}$ trivializes $\bar{\Phi}^{\gamma}$. Suppose first that $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is nontrivial. Fix $\vec{\eta}\in(\lim(\mu))^{n}$, and let $b\in\gamma^{n}$ be such that $c_{b(i)}=\{\delta^{\gamma}_{\xi}\mid\xi<\eta_{i}\}$ for all $i<n$. Unraveling the definitions and constructions yields the following facts.
@@
    For all $\xi<\eta_{0}$, we have $\psi^{*}_{b}(\delta^{\gamma}_{\xi})=\bar{\psi}^{\gamma}_{\vec{\eta}}(\xi)$.
@@
For $\vec{\zeta}\in(\lim(\mu))^{n-1}$, define $\nu_{\vec{\zeta}}:\zeta_{0}\rightarrow\mathbb{Z}$ as follows. Let $\vec{g}_{\vec{\zeta}}=\langle g_{\zeta_{i}}\mid i<n-1\rangle$, and let $\nu_{\vec{\zeta}}(\xi)=\psi_{\vec{g}_{\vec{\zeta}}{}^{\frown}\langle g\rangle}(\xi,\delta^{\gamma}_{\xi})$ for all $\xi<\zeta_{0}$. Then item (3) above implies that $\langle(-1)^{n}\nu_{\vec{\zeta}}\mid\vec{\zeta}\in(\lim(\mu))^{n-1}\rangle$ trivializes $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ restricted to $\lim(\mu)$, which is a contradiction.
Finally, assume that $\bar{\Sigma}^{\gamma}-\bar{\Psi}^{\gamma}$ is trivial, as witnessed by $\langle\rho_{\vec{\zeta}}\mid\vec{\zeta}\in\mu^{n-1}\rangle$. Fix $\vec{\eta}\in(\lim(\mu))^{n}$ and $b\in\gamma^{n}$ as in the previous paragraph. Item (1) of that paragraph is unchanged, and items (2) and (3) become
@@
For all $\vec{\zeta}\in(\lim(\mu))^{n-1}$, define $\nu_{\vec{\zeta}}:\zeta_{0}\rightarrow\mathbb{Z}$ as follows. Let $\vec{g}_{\vec{\zeta}}=\langle g_{\zeta_{i}}\mid i<n-1\rangle$, and let $\nu_{\vec{\zeta}}(\xi)=(-1)^{n-1}\rho_{\vec{\zeta}}(\xi)+\psi_{\vec{g}_{\vec{\zeta}}{}^{\frown}\langle g\rangle}(\xi,\delta^{\gamma}_{\xi})$ for all $\xi<\zeta_{0}$. Then the above calculation implies that $\langle\nu_{\vec{\zeta}}\mid\vec{\zeta}\in(\lim(\mu))^{n-1}\rangle$ trivializes $\langle\tau_{\eta}\mid\vec{\eta}\in(\lim(\mu))^{n}\rangle$, yielding the final contradiction. ∎
@@
The case of $n=0$ is the ZFC result Theorem 2.10(5). For higher $n$, the relevant conditions of Theorem 2.20 all hold in Gödel’s constructible universe $\mathrm{L}$. To see this, first note that $\mathsf{GCH}$ holds in $\mathrm{L}$ and therefore, by [37, Theorem 3.40], $S^{\aleph_{n+1}}_{\aleph_{n}}\in I[\aleph_{n+1}]$ for all $1\leq n<\omega$. By a theorem of Jensen [50], $\diamondsuit(S^{\aleph_{n+1}}_{\aleph_{n}})$ holds in $\mathrm{L}$. Finally, there exists a nontrivial $n$-coherent family $\langle\tau_{\vec{\eta}}:\bigwedge\vec{\eta}\rightarrow\mathbb{Z}\mid\vec{\eta}\in(\omega_{n})^{n}\rangle$ in $\mathrm{L}$ by [17, Cor. 3.27]. ∎